OpenStudy (anonymous):

Find an equation in general form for the straight line that passes thru the point (-1,4) and is perpendicular to the line 2x+3y=6

OpenStudy (anonymous):

slope of 2x +3y =6 is -2/3 Line perpendicular to this line will have slope 3/2 y = 3/2 x + b is the perpendicular line This line passes through (-1,4) Plug in x =-1 , y = 4 and solve for b 4 = 3/2(-1) +b 4 = -3/2 +b 4 +3/2 = b 11/2 =b The line is y = 3/2x + 11/2 2y = 3x +11 2y -3x = 11 is the required line